OpenStudy (anonymous):

Factor 18s-63 I'll give you a medal :)

OpenStudy (jdoe0001):

\(\bf 18s-63\implies (2\cdot 9)-(7\cdot 9)\)

OpenStudy (jdoe0001):

well.. \(\bf 18s-63\implies (2\cdot 9)s-(7\cdot 9)\)

OpenStudy (anonymous):

9(2s-7)

OpenStudy (anonymous):

Thank you do much^-^

OpenStudy (anonymous):

So*

OpenStudy (anonymous):

u r welcom